the perimeter of a rectangle is 480 ft and the sides are in a ratio of 5:1. what are the length and width of the two sides of the rectangle
Answer by macston(5194)

for width =W ,length=L=5W
perimeter=2L+2W=480 ft
so 2(5W) + 2W = 480
10W+2W=480
12W=480
W=40 ANSWER 1
L=5W
L=5(40)=200 ANSWER 2
Check:
Perimeter=L+W+L+W
480=200+40+200+40
480=480